OpenStudy (anonymous):

Simplify.square root of 10 times square root of 8

OpenStudy (tylerd):

\[\sqrt{10} \times \sqrt{8} = \sqrt{10 \times 8}\] first step

OpenStudy (tylerd):

what do you think we'd do next?

OpenStudy (anonymous):

no clue I'm terrible at radicals @TylerD

OpenStudy (tylerd):

well the \[\sqrt{10 \times 8}\] we can evalute to = \[\sqrt{80}\]

OpenStudy (tylerd):

when dealing with radicals we do something like this \[\sqrt{25} = \sqrt{5 \times 5}\]

OpenStudy (tylerd):

when there are 2 of the same number multiplied by eachother inside a square root, we move it outside, so in that case the square root of 25 becomes 5

OpenStudy (anonymous):

makes sense

OpenStudy (tylerd):

so for a larger number like 80 it takes a bit more work.

OpenStudy (anonymous):

Whats the biggest square you can find that fits into 80?

OpenStudy (tylerd):

however and easier way to do it

OpenStudy (tylerd):

10 x 8 is already simplified for us

OpenStudy (tylerd):

so rather then multiply to get 80 we could actually just break those numbers up

OpenStudy (anonymous):

wow you're getting a metal thanks soo much

OpenStudy (tylerd):

do you got it?

OpenStudy (anonymous):

yes :) @TylerD

OpenStudy (anonymous):

can you help with a couple of more? you don't have too

OpenStudy (tylerd):

ya i can help with one more

OpenStudy (anonymous):

thats not an option so i guess its none of the above

OpenStudy (anonymous):

square root of 18 4 square root of 5 square root of 80 None of the above

OpenStudy (tylerd):

well \[\sqrt{80} = \sqrt{2 \times 2 \times 2 \times 2 \times 5}\]

OpenStudy (tylerd):

so we move 2 2's out it becomes 4sqrt(5) my error

OpenStudy (anonymous):

For the original answer I got 4 times square root of 5

OpenStudy (tylerd):

yes you were right

OpenStudy (anonymous):

so none of the above is correct @TylerD

OpenStudy (anonymous):

B

OpenStudy (tylerd):

no 4sqrt5

OpenStudy (tylerd):

we have four 2's times 5 inside the square root so we have 2 pairs of 2's when you move something outside of the square root your multiplying

OpenStudy (tylerd):